Please help
(9%) Problem 8: A 0.26-kg aluminum bowl holding 0.65 kg of soup at 25.0°C is placed in a freezer. Specific heat (c) ubstance Solids Aluminum Concrete 900 840 387 840 2090 0.20 0.0924 0.20 Glass Ice (average Liquids Water 4186 1.000 Gases Steam (100C) 1520 (2020) 0.363(0.482) Ctheexpertta.com What is the final temperature, in degrees Celsius, if 377 kJ of energy is transferred from the bowl and soup, assuming the soups thermal properties are the same as that of water? Grade Summary Deductions 8% 92% coso tan0 acoso acotansinho 78 9 sino cotan asin atan) Attempts remaining (2% per attempt) detailed view I 4 5 6 tanh cotanh0 2% O Degrees O Radians 2% Hint I give upt Hints:-% deduction per hint. Hints remnanning:- Feedback:-deduction per feedback. Submission History Hints Feedback Totals Answer Tr 29.17 h-.29.17 Tr- 29 056 | Note: Feedback not accessed. 0% 12% 11% 12% 2% 0% = 152.58 -1526 0% Tr 163.5 2% 0% 12% Totals 0% 18%
Answer #1

the formula to get final temperature is

377\times 10^3 +900\times (T-25)\times 0.26 + 4186\times (T-25)\times 0.65=0

which will give us T = -102.6 C